# mimic-video: Video-Action Models for Generalizable Robot Control Beyond VLAs

### v1

@@ I Introduction
Recent progress in general-purpose robotic manipulation has been driven by the paradigm of imitation learning, particularly when combined with large-scale pretraining. State-of-the-art methods, often termed Vision-Language-Action models (VLAs), finetune large pretrained Vision-Language Models (VLMs) on robotics datasets [59, 26, 3, 24]. By leveraging semantic knowledge distilled from vast quantities of internet images and text, these models can follow natural language instructions, exhibit broad generalization to novel scenarios, and learn a wide range of skills with a single policy.
@@
However, this paradigm faces a fundamental limitation: the pretraining data, while massive in scale, is inherently static. Images and text lack the explicit, temporally-grounded information about dynamics and physical procedures that are crucial for complex manipulation. By default, a VLM has no incentive to learn intuitive physics or the step-by-step evolution of a scene during its vision-language pretraining phase. Consequently, the burden of learning physical dynamics (how objects move, deform, and interact) falls entirely on the post-training stage, where the model must infer these from scarce and expensive expert-teleoperated demonstrations. This heavy reliance on robot data creates a data-efficiency bottleneck that limits scalability. Furthermore, standard VLAs typically necessitate backpropagating action gradients through the entire vision-language backbone. Consequently, gradients from randomly initialized action adapters unfavorably interact with the pretrained weights, causing gradient interference that degrades the model’s semantic knowledge and language-following capabilities [14].
@@
In this work, we posit that the key to more sample-efficient and capable robot policies lies in leveraging a pretraining modality that inherently encodes dynamic, procedural information: video. Unlike static image-text pairs, internet-scale video data provides rich knowledge on “how things are done”, capturing the nuanced physics of interaction, how objects move, deform and react to forces. However, effectively harnessing this data remains a significant challenge. Prior approaches often attempt to retrofit human videos by pre-processing them to extract low-level pseudo-actions [47, 50, 25]. We argue that this strategy is fundamentally limited: the extraction process is prone to significant noise and artifacts, and it introduces a severe embodiment gap.
@@
To address these limitations, we propose a more direct paradigm: grounding robot policies directly within the latent representations of a generative, pretrained video model. We introduce mimic-video, a novel Video-Action Model (VAM) that unifies large-scale video prediction with robot control. Built upon a state-of-the-art video diffusion backbone, mimic-video functions by first synthesizing a visual plan: given an initial observation and language instruction, the video backbone partially generates a future trajectory within a compact, tokenized latent space. During video generation, we extract intermediate representations from the video model that serve as a plan conditioning a downstream action decoder, which operates as an Inverse Dynamics Model (IDM) to recover the corresponding low-level motor commands. Fundamentally, this architecture decouples the inherent multi-modality of long-horizon planning, now offloaded to the video backbone, from the downstream control task. This effectively frees the action decoder from modeling complex future distributions, allowing it to dedicate its entire capacity to the far simpler, unimodal and non-causal problem of inverse dynamics [34, 35].
@@
Our primary contribution is mimic-video, a novel generalist robot policy that integrates generative video pretraining with flow matching-based control, establishing a new class of methods we term Video-Action Models (VAMs). We evaluate our approach across a diverse suite of robotic embodiments—ranging from standard single-arm manipulation to bimanual dexterous tasks, demonstrating state-of-the-art results in both simulated benchmarks and challenging real-world environments. Our mimic-video model achieves this performance while improving sample efficiency by 10x and convergence speed by 2x compared to traditional VLA architectures.
@@ II Related Work
End-to-end imitation learning has become the dominant paradigm for training general-purpose robot manipulation policies, enabling robots to acquire complex skills directly from expert demonstrations. This approach, which maps raw sensory observations to actions, has benefited from advances in generative modeling in so far as models act as “data sponges”, able to absorb large and diverse pretraining datasets [10, 26, 59, 48, 12, 45, 4] to achieve downstream generalization in action generation.
While early approaches like ACT [57] used a VAE to model action chunks, the field has shifted toward iterative generative frameworks, popularized by Denoising Diffusion Models [21]. This class of methods, encompassing Diffusion Policy [8, 11] and the Flow Matching [31] decoders of $\pi$ [3, 24], has become the state-of-the-art. These generative approaches excel at modeling multi-modal distributions of expert actions and form the technical foundation for modern robot imitation learning policies, including our own action decoder.
@@
A major breakthrough in robot learning has been the advent of Vision-Language-Action (VLA) models, which are obtained by finetuning large, pretrained Vision-Language Models (VLMs) on robotics data. Models like RT-2 [59], OpenVLA [26], and the $\pi_{0}$$\pi_{0.5}$ series [3, 42, 24] leverage the vast semantic and visual knowledge embedded in their backbones from pretraining on internet-scale image-text data. This allows them to follow open-ended language instructions, understand abstract concepts, and generalize to novel objects and tasks in a zero-shot fashion. However, a fundamental limitation of VLAs is that the VLM backbones they make use of are only pretrained with static image data and language descriptions. They lack an inherent model of video dynamics, physics, or temporal progression, limiting their ability to reason about the physical consequences of actions. This critical knowledge must be learned from scratch from comparatively small and expensive robotics datasets.
@@
The utilization of video prediction for robotic control has a long-standing history, primarily motivated by the potential to enable planning through visual foresight. Early works, such as the seminal approaches by Oh et al. [39], Watter et al. [51], Fragkiadaki et al. [19], Finn et al. [18], Finn and Levine [17], demonstrated how video prediction could enhance physical interaction. As generative models have matured to produce high-definition, coherent long-form content [40, 53], recent works have explored diverse integrations of video generation with policy learning.
Some recent examples include: Dreamitate [29], finetuning a pretrained video generator and then explicitly tracking end effector-mounted tools in videos to obtain actions; Video Language Planning [16], Generative Predictive Control (GPC) [43] and V-JEPA 2 [1], learning predictive world models from which goal-conditioned action plans can be extracted; Unified World Models (UWM) [58] and Video Policy [30], in which models are co-trained from scratch in a purely end-to-end fashion to jointly model actions and video dynamics; UniPi [15], learning a video generator and separately a pixel-to-action Inverse Dynamics Model (IDM); and LAPA [54], first extracting latent actions from video, and then aligning a VLM to these action representations and training a VLA policy.
Our proposed approach departs from prior work by directly grounding control in the rich latent priors of internet-scale video models, rather than training from scratch or relying on pixel-level reconstruction. By conditioning a lightweight inverse dynamics model on intermediate, noisy latent states, we bypass the computational cost of full video generation and the brittleness of heuristic tracking. This enables a scalable, end-to-end framework that effectively transfers broad physical understanding to downstream manipulation tasks significantly reducing the reliance on expensive, large-scale robotic demonstrations.
Figure 0: Policy performance scales with video generation quality. We compare success rates when conditioning our action decoder on different visual inputs: video latents generated by either predictions or ground-truth (expert) video for both features from a standard pretrained video model (gray), as well as a video model finetuned on video data from the robot dataset (red). The near-perfect performance with ground truth inputs confirms that control effectively reduces to visual prediction, implying policy performance scales directly with video model quality.
@@ III Case Study: How does Video Generation Quality Affect Robot Policy Performance?
In this work, we argue that the internal representations arising in pretrained video model backbones are better suited for downstream robot learning compared to those in the VLM backbones commonly used in current state of the art VLAs. Intuitively, video models jointly model images, physical dynamics, as well as visual action plans. A policy trained on such representations effectively reduces the role of the action decoder to a simple translator, mapping visual action plans into a low-dimensional robot action trajectory. If this hypothesis holds, the bulk of learning in Video-Action Models falls on the large-scale video pretraining and finetuning phases, while training the action decoder (the step requiring expensive, high-quality robot teleoperation data) becomes lightweight and data efficient.
@@ IV Video-Action Models
We introduce mimic-video, a generative Video-Action Model (VAM) capable of modeling the joint distribution of video and robot actions. Our architecture couples two Conditional Flow Matching (CFM) models: a pretrained, language-conditioned video backbone, and a lightweight action decoder that functions as an Inverse Dynamics Model (IDM) by conditioning on the video model’s latent representations.
@@ IV-A Preliminaries: Flow Matching
Both the video and action prediction components are trained using the Flow Matching framework [31] to model a data distribution $p_{0}(x^{0})$ by constructing a Continuous Normalizing Flow [6]. We use the optimal transport path
@@
which interpolates between clean data $x^{0}$ (at $\tau=0$) and Gaussian noise $x^{1}=\varepsilon\sim\mathcal{N}(0,I)$ (at $\tau=1$) to define the conditional probability path $p_{\tau}\left(x^{\tau}\mid x^{0}\right)$. The model parameterizes an estimator $v_{\theta}$ to the intractable marginal generating vector field
@@
Inference is performed by integrating the learned field $v$ from $\tau=1$ to $\tau=0$ to recover $\hat{x}^{0}\sim p_{0}$:
$\hat{x}^{0}=\epsilon+\int_{1}^{0}v_{\theta}(\hat{x}^{\tau},\tau)d\tau$ (3)
Critically, this continuous time parameter $\tau$ allows us to define partial denoising (stopping at intermediate $\tau_{v}>0$), which is central to our method.
Figure 1: mimic-video architecture: we instantiate our framework with a pretrained video generation backbone (Cosmos-Predict2 [38, 37]), which provides rich physical dynamics priors learned from large-scale video data. We adapt this model for control via a partial denoising strategy, where the video backbone tracks flow to an intermediate noise level $\tau_{v}$ to extract latent visual plans. These representations condition a smaller action decoder, which processes proprioceptive states and predicts action trajectories. The video and action components operate on independent flow schedules ($\tau_{v}$ and $\tau_{a}$), allowing us to design the learning problem optimally for each modality.
@@ IV-B Architecture Formulation
Formally, we aim to learn a generalist robot policy $\pi(\mathbf{A}_{t}\mid\mathbf{o}_{t},\,l)$ that predicts a sequence of actions $\mathbf{A}_{t}=[\mathbf{a}_{t},\dots,\mathbf{a}_{t+H-1}]$ given observations consisting of multiple RGB images $\mathbf{I}_{t}$, a language instruction $l$ and the robot’s proprioceptive state $\mathbf{q}_{t}$, such that $\mathbf{o}_{t}=[\mathbf{I}_{t}^{(1)},\dots,\mathbf{I}_{t}^{(N)},l,\mathbf{q}_{t}]$.
@@ IV-D Action Decoder
The action decoder is instantiated as a DiT that encodes the robot’s proprioceptive state $\mathbf{q}_{t}$ and a sequence of $\mathbf{A}_{t}$ future robot actions through two separate MLP networks and concatenates them to form the action decoder’s sequence dimension. We use learned absolute positional encodings to add temporal information to each token. During training, we randomly replace the soft token encoding the proprioceptive state with a learned mask token to prevent overfitting on the low dimensional observation. Each action decoder layer consists of (1) cross-attention to intermediate video model representations $\mathbf{h}_{\tau_{v}}$, (2) self-attention over the action sequence, and (3) a two-layer MLP. Each component is bypassed by a residual path and each component’s output is modulated via AdaLN [41], where the input to the AdaLN projections is a low-rank bilinear-affine encoding of both video and action flow times $\tau_{v}$ and $\tau_{a}$.
@@ IV-E Action Sampling
To enable real-time control, we formulate inference as efficient sampling from the marginal action policy. Although mimic-video is in principle capable of sampling from the joint video-action distribution, we can sample from the marginal action distribution more efficiently by bypassing the computational cost of full video reconstruction. We therefore propose a partial denoising strategy that extracts semantic features from intermediate flow states without resolving fine-grained pixel details. Our inference action sampling procedure is described in Algorithm 1. Given image observations $\mathbf{o}_{t}$, we integrate the video flow field from Gaussian noise to an intermediate flow time $\tau_{v}$ (see Eq. 3). This yields a partially denoised latent state $\mathbf{z}^{\tau_{v}}_{\text{future}}$ that retains sufficient structural information to guide the policy. We process this state with the first $k$ layers of the video model and pass the resulting activations as conditioning information to the action decoder. The action decoder then performs a full denoising procedure to produce a chunk of robot actions $\mathbf{A}^{0}_{t}$.
@@
At inference time, $\tau_{v}$ is a free hyperparameter. Its optimal value is task-dependent, but we show in Sec. V-C empirically that it is generally close to $1$ (high noise). In the special case of $\tau_{v}=1$, a single forward pass of the computationally intensive video backbone is sufficient to generate a chunk of actions (line 3 in Algorithm 1 becomes redundant), facilitating real-time inference in our experiments. We find that $\tau_{v}=1$ is a good default value that balances policy performance and inference speed. See Sec. -D for a discussion on the motivation behind “noisy” video conditioning.

### v2

@@ I Introduction
Building on the capabilities of pretrained Vision-Language Models (VLMs), Vision-Language-Action Models (VLAs) transfer semantic knowledge acquired through Internet-scale vision-language pretraining to the physical domain. By finetuning a VLM on diverse robot action data, VLAs learn generalist, natural language-conditioned robot manipulation policies that combine a wide range of skills and exhibit impressive generalization to unseen instructions, objects, and environments [60, 26, 3, 25].
@@
However, this paradigm faces a fundamental limitation: the pretraining data, while massive in scale, is inherently static. Images and text lack explicit, temporally-grounded information about dynamics and physical procedures that are crucial for complex manipulation. Consequently, the burden of learning physical dynamics (how objects move, deform and interact) falls entirely on the post-training stage, where the model must infer these from scarce and expensive expert-teleoperated demonstrations. This heavy reliance on robot data creates a data-efficiency bottleneck that limits scalability. While prior works have explored augmenting VLA training with auxiliary video-derived signals such as language plans, affordances, or keypoints [55, 19, 24, 25], reducing dense video into such sparse representations creates an information bottleneck, failing to capture fine-grained dynamics.
@@
In this work, we posit that the key to more sample-efficient and capable robot policies lies in leveraging a pretraining modality that inherently encodes dynamic, procedural information: video. Unlike static image-text pairs, internet-scale video data provides rich knowledge on “how things are done”, capturing the nuanced physics of interaction, how objects move, deform and react to forces. However, effectively harnessing this data remains a significant challenge. Prior approaches leveraging pretrained video models typically learn the joint distribution over video and actions, factorized such that the predicted actions are conditioned on synthesized future frames [30, 29, 15]. However, recovering the policy typically requires fully generating these future frames, necessitating prohibitive video synthesis at every control step.
@@
To address these limitations, we propose a more direct paradigm: grounding robot policies directly in the latent representations of a generative, pretrained video model. We introduce mimic-video, a novel Video-Action Model (VAM) that unifies video modeling with robot control. Built upon a state-of-the-art video diffusion backbone, mimic-video functions by first synthesizing a visual plan: given an initial observation and language instruction, the video backbone predicts a future trajectory within a compact latent space. Rather than requiring full or even partial video generation, we extract intermediate video model representations to condition a downstream action decoder. This decoder operates as an Inverse Dynamics Model (IDM) to recover low-level motor commands. This formulation allows the video backbone to remain frozen, eliminating the need to train it on scarce robot action data. Fundamentally, this architecture decouples the inherent multi-modality of long-horizon planning, now offloaded to the video backbone, from the downstream control task. This effectively frees the action decoder from modeling complex future distributions, allowing it to dedicate its entire capacity to the far simpler, unimodal and non-causal problem of inverse dynamics [34, 35].
@@
Our primary contribution is mimic-video, a novel generalist robot policy that integrates generative video pretraining with flow matching-based control, establishing a new class of methods we term Video-Action Models (VAMs). We evaluate our approach across a diverse suite of robotic embodiments ranging from standard single-arm manipulation to bimanual dexterous tasks, demonstrating state-of-the-art results in both simulated benchmarks and challenging real-world environments. Our mimic-video model achieves this performance while improving sample efficiency by 10x and convergence speed by 2x compared to traditional VLA architectures.
@@ II Related Work
End-to-end imitation learning has become the dominant paradigm for training general-purpose robot manipulation policies, enabling robots to acquire complex skills directly from expert demonstrations. This approach, which maps raw sensory observations to actions, has benefited from advances in generative modeling in so far as models act as “data sponges”, able to absorb large and diverse pretraining datasets [9, 26, 60, 11, 45, 4] to achieve downstream generalization in action generation.
While early approaches like ACT [57] used a VAE to model action chunks, the field has shifted toward iterative generative frameworks, popularized by Denoising Diffusion Probabilistic Models [22]. This class of methods, encompassing Diffusion Policy [8, 10] and the Flow Matching [31] decoders of the $\pi_{0}$$\pi_{0.5}$ series [3, 25], has become the state-of-the-art. These generative approaches excel at modeling multi-modal distributions of expert actions and form the technical foundation for modern robot imitation learning policies, including our own action decoder.
@@
A major breakthrough in robot learning has been the paradigm of Vision-Language-Action (VLA) models, which are obtained by finetuning large, pretrained Vision-Language Models (VLMs) on robotics data. Models like RT-2 [60], OpenVLA [26], and the $\pi_{0}$$\pi_{0.5}$ series [3, 42, 25] leverage the vast semantic knowledge embedded in their backbones from pretraining on internet-scale image-text data. This allows them to follow open-ended language instructions, understand abstract concepts, and generalize to novel objects, environments, and tasks in a zero-shot fashion. However, a fundamental limitation of VLAs is that the VLM backbones they make use of are only pretrained with static vision and language data. They lack an inherent model of video dynamics, physics, or temporal progression, limiting their ability to reason about the physical consequences of actions. This critical knowledge must be learned from scratch from comparatively small and expensive robotics datasets.
@@
The utilization of video prediction for robotic control has a long-standing history, primarily motivated by the potential to enable planning through visual foresight. Early works, such as the seminal approaches by Oh et al. [39], Watter et al. [51], Fragkiadaki et al. [18], Finn et al. [17], Finn and Levine [16], demonstrated how video prediction could enhance physical interaction. As generative models have matured to produce high-definition, coherent long-form content [40, 53], recent works have explored diverse integrations of video generation with policy learning. The use of action-conditioned video models (world models) for policy learning has recently seen significant adoption. World models can help select more optimal action sequences at runtime by "imagining" their outcome [1, 43], or be used as learned simulators for evaluation and DAGGER-like [47] approaches [20, 48]. In this work, we consider non action-conditioned video models. One line of work fully generates pixel-space future video and obtains actions either via non-parametric methods such as tracking a custom end effector-mounted tool [29], or learned pixel-based Inverse Dynamics Models [14, 15]. CoT-VLA [56] uses a pretrained VLM capable of generating images to generate a subgoal image and actions in one autoregressive sequence. Another line of work learns to model video during training without predicting video in action sampling; Unified World Models [59] learns a model from scratch that can flexibly function as a policy, a video prediction model, or a forward or inverse dynamics model and LAPA [54] finetunes a VLM to predict “latent actions” (an encoding of the difference between the current and a future image), re-training only the output layer to predict actions in a subsequent training stage. FLARE [58] aligns intermediate VLA representations with future vision-language embeddings, implicitly modeling video and actions jointly. Similarly to our approach, Video Policy [30] explicitly models the joint video-action distribution and conditions a policy model on intermediate video model representations, but crucially does not allow for efficient sampling of the marginal action distribution.
Our proposed approach departs from most prior work by directly grounding control in the rich latent priors of internet-scale video models, rather than training from scratch or relying on pixel-level reconstruction. Additionally, by conditioning a lightweight inverse dynamics model on intermediate, noisy latent states, we bypass the computational cost of full video generation and the brittleness of heuristic tracking. This enables a scalable, end-to-end framework that effectively transfers broad physical understanding to downstream manipulation tasks, significantly reducing the reliance on expensive, large-scale robotic demonstrations.
Figure 0: We compare success rates when conditioning our action decoder on different visual inputs: video latents generated by either predictions or ground-truth (expert) video for both features from a standard pretrained video model (gray), as well as a video model finetuned on video data from the robot dataset (orange). The near-perfect performance with ground truth inputs confirms that control effectively reduces to visual prediction, implying policy performance scales directly with video model quality.
@@ III Case Study: How does Video Generation Quality Affect Robot Policy Performance?
In this work, we argue that the internal representations arising in pretrained video model backbones are better suited for downstream robot learning compared to those in the VLM backbones commonly used in current state of the art VLAs. Intuitively, video models jointly model images, physical dynamics, alongside visual action plans. A policy trained on such representations effectively reduces the role of the action decoder to a simple translator, mapping visual action plans into a low-dimensional robot action trajectory. If this hypothesis holds, the bulk of learning in Video-Action Models falls on the large-scale video pretraining and finetuning phases, while training the action decoder (the step requiring expensive, high-quality robot teleoperation data) becomes lightweight and data efficient.
@@ IV Video-Action Models
We introduce mimic-video, a generative Video-Action Model (VAM) capable of modeling the joint distribution of video and robot actions. Our architecture couples two Conditional Flow Matching (CFM) models: a pretrained, language-conditioned video backbone and a lightweight action decoder that functions as an Inverse Dynamics Model (IDM) by conditioning on the video model’s latent representations.
Figure 1: mimic-video architecture: we instantiate our framework with a pretrained video generation backbone (Cosmos-Predict2 [38, 37]), which provides rich physical dynamics priors learned from large-scale video data. We adapt this model for control via a partial denoising strategy, where the video backbone follows the flow to an intermediate flow time $\tau_{v}$ to extract latent visual plans. These representations condition a smaller action decoder, which processes proprioceptive states and predicts action trajectories. The video and action components operate on independent flow schedules ($\tau_{v}$ and $\tau_{a}$), allowing us to design the learning problem separately for each modality.
@@ IV-A Preliminaries: Flow Matching
Both the video and action prediction components are trained using the Flow Matching framework [31] to model a data distribution $p_{0}(x^{0})$ by constructing a Continuous Normalizing Flow [6]. We use the conditional optimal transport path
@@
which interpolates between clean data $x^{0}$ (at $\tau=0$) and Gaussian noise $\varepsilon\sim\mathcal{N}(0,I)$ (at $\tau=1$) to define the conditional probability path $p_{\tau}\left(x^{\tau}\mid x^{0}\right)$. The model parameterizes an estimator $v_{\theta}$ to the intractable marginal generating vector field
@@
Inference is performed by integrating the learned field $v_{\theta}$ from $\tau=1$ to $\tau=0$ to recover $\hat{x}^{0}\sim p_{0}$:
$\hat{x}^{0}=\varepsilon+\int_{1}^{0}v_{\theta}(\hat{x}^{\tau},\tau)d\tau$ (3)
Critically, this continuous time parameter $\tau$ allows us to define partial denoising (stopping at intermediate $\tau>0$), which is central to our method.
@@ IV-B Architecture Formulation
Formally, we aim to learn a generalist robot policy $\pi(\mathbf{A}_{t}\mid\mathbf{o}_{t},\,l)$ that predicts a sequence of actions $\mathbf{A}_{t}=[\mathbf{a}_{t},\dots,\mathbf{a}_{t+H_{a}-1}]$ given observations consisting of multiple RGB images $\mathbf{I}_{t^{\prime}}$, a language instruction $l$ and the robot’s proprioceptive state $\mathbf{q}_{t}$, such that $\mathbf{o}_{t}=[\mathbf{I}_{t-H_{o}+1},\dots,\mathbf{I}_{t},l,\mathbf{q}_{t}]$.
@@ IV-D Action Decoder
The action decoder is instantiated as a DiT that encodes the robot’s proprioceptive state $\mathbf{q}_{t}$ and a sequence of $\mathbf{A}_{t}$ future robot actions through two separate MLP networks and concatenates them to form the action decoder’s sequence dimension. We use learned absolute positional encodings to add temporal information to each token. During training, we randomly replace the soft token encoding the proprioceptive state with a learned mask token to prevent overfitting on the low dimensional observation. Each action decoder layer consists of (1) cross-attention to intermediate video model representations $\mathbf{h}^{\tau_{v}}$, (2) self-attention over the action sequence, and (3) a two-layer MLP. Each component is bypassed by a residual path and each component’s output is modulated via AdaLN [41], where the input to the AdaLN projections is a low-rank bilinear-affine encoding of both video and action flow times $\tau_{v}$ and $\tau_{a}$.
@@ IV-E Action Sampling
To enable real-time control, we formulate inference as efficient sampling from the marginal action policy. Although mimic-video is in principle capable of sampling from the joint video-action distribution (see Fig. 2 for an example), we can sample from the marginal action distribution more efficiently by bypassing the computational cost of full video reconstruction. We therefore propose a partial denoising strategy that extracts semantic features from intermediate flow states without resolving fine-grained pixel details. Our inference action sampling procedure is described in Algorithm 1. Given image observations $\mathbf{o}_{t}$, we integrate the video flow field from Gaussian noise to an intermediate flow time $\tau_{v}$ (see Eq. 3). This yields a partially denoised latent state $\mathbf{z}^{\tau_{v}}_{\text{future}}$ that retains sufficient structural information to guide the policy. We process this state with the first $k$ layers of the video model and pass the resulting activations as conditioning information to the action decoder. The action decoder then performs a full denoising procedure to produce a chunk of robot actions $\mathbf{A}^{0}_{t}$.
@@
At inference time, $\tau_{v}$ is a free hyperparameter. Its optimal value is task-dependent, but we show in Sec. V-C empirically that it is generally close to $1$ (high noise). In the special case of $\tau_{v}=1$, a single forward pass of the computationally intensive video backbone is sufficient to generate a chunk of actions (line 3 in Algorithm 1 becomes redundant), facilitating real-time inference in our experiments. We find that $\tau_{v}=1$ is a good default value that balances policy performance and inference speed. See Sec. -E for a discussion on the motivation behind “noisy” video conditioning.